# revtex4-1Repair the float Journal of UAP Studies        Society for UAP Studies

### v2

@@ 3 Results
Figure 2 presents all the sim results, divided by annual probability and cumulative (integrated) probability vs. time, and also separated by start year, 2008 or 2024, with cases having the latter as the input essentially taking the current state of affairs (lack of proof on the web) as a (Bayesian) prior. In those latter situations, over-optimistic postulates can artificially generate enormous immediate peaks in probability, for 2024–2025. Using our “central” postulates, the predictions are 2038 $\pm$ 24 (2008 sim start) and 2049 $\pm$ 23 (2024 start) for the year of the initial but “incontrovertible” evidence being shared on the internet (assuming survival of strict checks of fakery). The uncertainties $i.e.$ errors quoted here are non-Gaussian – they are simply raw standard deviations $\sigma$ of asymmetric data. The mean expectations are not the same as the median, RMS (root-mean-square) or mode (peak in probability), with that last parameter having no value higher than 2068 even when others extend into the 22nd century, due $e.g.$ to a global-pop peak.
Figure 2: Odds of catastrophic disclosure by year at left (a,c) on log axes, and summed at right (b,d), beginning in 2008 (top), and 2024 (bot). Pessimistic refers to all inputs “low” (Fig. 1), realistic medium, optimistic high. Roswell means 0.01 crash/year, Grusch 0.1/yr, and Randle 1, but * indicates he thinks that is far too high. Some stats are tabulated at the right.
@@ 4 Discussion and Conclusion
If NHI are real then the correct question to ask is not IF disclosure can be forced, but WHEN. Figure 2 right shows the probability always asymptoting to 100% eventually – when inputting the defaults, the chances (black curves) are 14–42% by 2027 and 39–59% by 2036. (Two values are quoted for each example year because of the differing beginning years.) This is because of the remarkable interconnectedness of the modern world, with cellular phones in individuals’ pockets and bags nearly everywhere even in developing nations, even when one makes the conservative claim of only a fixed 54% maximum for the percentage of human beings with camera-enabled smartphones. We did not account for people who have multiple phones and subscriptions, but this could be the cause of the significant tension between the sources we cited earlier (50 as opposed to about 75%), and this is why it is so important that divergent scenarios were studied. A further refinement to our study that readers can implement on their own since the code has been provided at the end is a mixing of low, medium, and high inputs instead of their simultaneous railing, which was done here for simplicity. Future work building upon this paper by this author and/or others could also account for daylight versus night-time hours, terrain type, attention spans, and other such factors, but the analyses presented herein represent a first stab, the first in the UAP literature as far as the author knows, at least in terms of serious (journal) publications.
@@
An unspoken assumption baked into every analysis within this manuscript is that we are referring to unmistakable, clear photos of crashes on land only, potentially less ephemeral than UFOs in the sky, including if not especially over waters Sanderson and Childress (2005); Dennett (2018), but lands account for only 29.2% of the surface area of Earth. Therefore, $\langle r\rangle$ = 0.1/year (our middle-benchmark crash rate) becomes $>$ 0.3/year for the entire globe, with $>$ 0.2/year for only the water-covered regions. If one equates any NHI with ETs, and takes the estimated time frame of the first discovery of Earth by a space-faring race capable of interstellar travel as $\sim$1 million years ago Knuth (2024), then this implies over 200,000 defunct extraterrestrial vessels sitting on the bottom of the Pacific Ocean, and other oceans, and seas, lakes, and rivers. This is a very surprising number which motivates further searches like that described in Loeb et al. (2024), and this might even be an underestimate – one species may inform others of Earth’s and of humanity’s presence, possibly leading to a time-dependent $r$-value (increasing?) driven by the scientific curiosity of others.
@@
Even though this work may read like it is a warning to government officials with knowledge of NHI or ET (again– should they really exist) that their time grows short to maintain full control of the narrative, it should not just sound alarm bells. The mathematics is neutral, in the sense that one could also capitalize on the procedures contained here to debunk NHI and UAP crash retrieval claims, as more years without disclosure pass by. The statistical techniques employed in this article can and should also be applied to cryptids, orbs, Earth lights, ball lightning, and any sort of “paranormal” claim that presently can boast of no proof that has convinced the majority of the scientific community, nor the public at large. The author was inspired by working on dark matter particles, a mainstream scientific topic with an enormous number of indirect i.e. observational clues coming from cosmology and astrophysics, but no conclusive, direct evidence in the laboratory as yet – one can set limits on the probability of their interaction with normal matter composed of atoms Aalbers et al. (2023). A scientist can set a limit on the rate of occurrence of any rare phenomenon using similar techniques, but without necessarily ruling it out completely. Not all exotic claims are false, as our history has repeatedly shown (atoms, germs, continental drift, meteors, air/space flight, nuclear power/bombs, relativity,…)
@@
That being said, it would be a mistake to state all initially “crazy” ideas have eventually been proven true within the history of science, as such a claim would be far from accurate. However, subjective experiences do not constitute unambiguous final evidence in the physical sciences, making the raw data sets of well-calibrated scientific instruments absolutely critical to possess, not just information from the “human sensor,” witnesses with potentially faulty memories (adequate for legal systems and useful in the humanities and the social sciences, but insufficient in a physical science). While a smartphone is not such an instrument, it can collect what Prof. Garry Nolan referred to at the Sol Foundation inaugural meeting as “pre-data,” a step above anecdotal non-data at least, which can still be used for corroborating scientific data, and justify choices of sensors. But, without funding and publications in mainstream journals, progress will still be difficult. A “smoking gun” phone pic could precipitate increases in both of those, but would still not be a substitute for the study of crash parts in person, to look for evidence of NHI technology, such as advanced unknown alloys, or isotopic concentrations inconsistent with our solar system Nolan et al. (2022). The latter could be observed using mass spectrometry, or non-destructively using NAA (neutron activation analysis) Laine et al. (2023). While an initially agnostic approach to data-taking is important Szydagis et al. (2023) scientists must be permitted, sans fear of reputation loss, to entertain “exotic” hypotheses (NHI, ET) and consider if data do favor them Villarroel (2023).
Acknowledgements
@@
The author acknowledges Prof. Eric Davis as well as David Grusch for inspiring this paper, and Mick West for his skepticism, and xkcd 1235. He thanks Prof. Kevin Knuth for first getting him interested in UAPs, and Gary Voorhis, leader of UAPx. He also thanks Prof. Brenda Denzler and Prof. Beatriz Villarroel for crucial criticisms of Version 1.
@@ Appendix A Supporting Code
The C++ file catDisc.cpp is attached for the benefit of the technically-minded reader with knowledge of computer programming. It can, for instance, be easily compiled on Unix with the terminal command “g++ -Ofast catDisc.cpp -o catDisc.out” assuming that one has the g++ compiler installed (the -Ofast optimization flag is optional).
@@ Appendix B Supporting Plots
Figure 3: The year of accidental disclosure with uncertainty as a function of $\langle r\rangle$, with all other numerical assumptions held fixed at their “medium” levels, for human population growth, smartphone penetration into the population, and visibility radius (in black: $R$ is varied in gray). The inset is a zoom-in for 10–100 crashes/century. $r>20$ is disfavored by sims at the Powell Radius (0.15 km). Discrete sim results are well fit continuously using y = 2008 + $m_{1}/r^{m_{2}}+m_{3}/r^{m_{4}}$, with both powers floating point.

### v3

@@ 3 Results
Figure 2 presents all the sim results, divided by annual probability and cumulative (integrated) probability vs. time, and also separated by start year, 2008 or 2024, with cases having the latter as the input essentially taking the current state of affairs (lack of proof on the web) as a (Bayesian) prior. In those latter situations, over-optimistic postulates can artificially generate enormous immediate peaks in probability, for 2024–2025. Using our “central” postulates, the predictions are 2038 $\pm$ 24 (2008 sim start) and 2049 $\pm$ 23 (2024 start) for the year of the initial but “incontrovertible” evidence being shared on the internet, assuming survival of strict checks of AI fakery. The errors (uncertainties) quoted here are non-Gaussian – they are simply raw standard deviations $\sigma$ of asymmetric data. The mean expectations are not the same as the median, RMS (root-mean-square) or mode (peak in probability), with that last parameter having no value higher than 2068 even when others extend into the 22nd century, due $e.g.$ to a global-population peak.
Figure 2: Odds of catastrophic disclosure by year at left (a,c) on log axes, and summed at right (b,d), beginning in 2008 (a,b), and 2024 (c,d). Pessimistic refers to all inputs “low” (Fig. 1), realistic medium, optimistic high. Roswell means 0.01 crash/year, Grusch 0.1/yr, and Randle 1, but * indicates Randle thinks that is far too high. Some statistics are tabulated at the right.
@@ 4 Discussion and Conclusion
If NHI are real then the correct question to ask is not IF disclosure can be forced, but WHEN. Figure 2 right shows the probability always asymptoting to 100% eventually – when inputting the defaults, the chances (black curves) are 14–42% by 2027 and 39–59% by 2036. (Two values are quoted for each example year because of the differing beginning years.) This is because of the remarkable interconnectedness of the modern world, with cellular phones in individuals’ pockets and bags nearly everywhere, even in developing nations, even when one makes the conservative claim of only a fixed 54% maximum for the percentage of human beings utilizing camera-enabled smartphones. We have not accounted for people who have multiple phones and subscriptions, but this could be the cause of the significant tensions between the sources we cited earlier ($\sim$50 as opposed to about 75%). Because of the contradictions, it is important that distinct scenarios were studied. A further refinement to our study which readers can implement on their own, as the code has been provided with the paper, is the mixture of low, medium, and high inputs in place of an artificial correlation. It was done here for simplicity. Future work building on this paper by this author and/or others could also account for daylight versus night-time hours, terrain type, attention spans, and other such factors Medina et al. (2023), but these analyses presented here are a first stab in the literature as far as the author knows, in terms of a journal publication.
@@
An unspoken assumption baked into every analysis within this manuscript is that we are referring to unmistakable, clear photos of crashes on land only, potentially less ephemeral than UAP in the sky, including if not especially over water Sanderson and Childress (2005); Dennett (2018). The land accounts for only 29.2% of the surface area of Earth. Therefore, $\langle r\rangle$ = 0.1/year (our middle-benchmark crash rate) becomes $>$ 0.3/year for the entire globe, with $>$ 0.2/year for only the water-covered regions. If one equates all NHI with aliens and takes the estimated time frame of the first discovery of Earth by one space-faring race capable of interstellar travel as $\sim$1 million years ago Knuth (2024), then this implies over 200,000 defunct (extraterrestrial) vessels sitting on the bottom of oceans, lakes, and rivers – if there was no change in technology over time. This is a surprising number motivating further searches like that described in Loeb et al. (2024), and this could even be an underestimate: one species may inform others of Earth and humanity’s presence and progress, possibly leading to a time-dependent $r$-value (increasing?) driven by others’ scientific curiosity.
@@
Although this work may read as if it is a warning to government officials with solid knowledge of some NHI (again– should they really exist) that their time grows short to maintain full control of the narrative, it should not just sound alarm bells. The mathematics is neutral, in the sense that one could also capitalize on the procedures contained here to debunk NHI and UAP crash retrieval claims, as more years without disclosure pass by. The statistical techniques employed in this article can and should also be applied to cryptids, orbs, Earth lights, ball lightning, and any sort of “paranormal” claim that presently can boast of no proof that has convinced the majority of the scientific community, nor the public at large. Using the dark matter as an example, a mainstream scientific topic with an enormous number of indirect i.e. observational clues coming from cosmology and astrophysics, but no conclusive, direct evidence in the laboratory as yet, we note one can set limits on the probability of the interaction of new particles with normal matter composed of atoms Aalbers et al. (2023). A scientist can set a limit on the rate of occurrence of any rare phenomenon with similar techniques, but without necessarily ruling it out completely. Not all exotic claims are false, as our history has repeatedly shown (atoms, germs, continental drift, meteors, air/space flight, nuclear power/bombs, relativity,…)
That being said, it would be a mistake to state that all initially “crazy” ideas have eventually been proven correct in the history of science, as such a claim would be far from accurate. However, subjective experiences do not constitute unambiguous final evidence in the physical sciences, making the raw data sets of well-calibrated scientific instruments absolutely critical to possess, not just information from the “human sensor,” witnesses with potentially faulty memories (adequate for legal systems and useful in the humanities and the social sciences, but insufficient in a physical science).
A central argument herein is that a phone may generate adequate evidence of an anomaly. Yet it is not a scientific instrument, so that is a major weakness we recognize in our own argument. While evidence from a single camera may convince the general public, it is not likely to convince most academics who continue to be unmoved by the existent plethora of UAP imagery. Having a video instead of photo(s) of a nearby crash on the ground, with a witness moving around and achieving different viewing angles, minimizes the possibility of an AI-generated hoax as a solution at the very least, especially if signs with letters and human fingers are also seen in the background Mirjalili (2023).
@@
Moreover, a smartphone camera can collect what Prof. Garry Nolan referred to at the Sol Foundation 2023 inaugural meeting as “pre-data,” a step above anecdotal non-data at least, which can still be used for corroborating scientific data, and justify choices of sensors. But, without funding and publications in mainstream journals, progress will still be difficult. A “smoking gun” phone video could precipitate increases in both of those, but would probably not be a substitute for the study of crash parts in person, to look for evidence of NHI technology, such as advanced unknown alloys, or isotopic concentrations inconsistent with our solar system Nolan et al. (2022). The latter could be observed using mass spectrometry, or non-destructively with NAA (neutron activation analysis) Laine et al. (2023). An initially agnostic approach to data-taking is good Szydagis et al. (2023) but scientists must be allowed, sans fear of reputation loss, to entertain exotic hypotheses (like NHI), considering if data favor them or not Villarroel and Krisciunas (2024).
Acknowledgments
@@
The author acknowledges Prof. Eric Davis as well as David Grusch for inspiring this paper, and Mick West for his skepticism, and xkcd 1235. He thanks Prof. Kevin Knuth for first getting him interested in UAPs, and Gary Voorhis, leader of UAPx. He thanks Profs. Brenda Denzler, Beatriz Villarroel, and Doug Buettner for key criticisms of drafts.
@@ Appendix A Supporting Code
The C++ file catDisc.cpp is part of a paper download for the benefit of the technically-minded reader with knowledge of computer programming. It can, for instance, be easily compiled on Unix with the terminal command “g++ -Ofast catDisc.cpp -o catDisc.out” assuming one has the g++ compiler installed (the -Ofast optimization flag is optional).
@@ Appendix B Supporting Plots
Figure 3: The year of accidental disclosure with uncertainty as a function of $\langle r\rangle$, with all other numerical assumptions held fixed, at their “medium” levels: for human population growth, smartphone penetration into the population, and visibility radius (in black: $R$ is varied in gray). The inset is a zoom-in for 10–100 crashes/century. $r>20$ is disfavored by sims at the Powell Radius (0.15 km). Discrete sim results are well fit continuously using y = 2008 + $m_{1}/r^{m_{2}}+m_{3}/r^{m_{4}}$, with both powers floating point. Smaller radii allow for the possibility of physical evidence retrieval by the crash witnesses, and may make hoaxing much harder, but the assumption of a single eyewitness was maintained: it is changeable to any number, however, in the code that is provided.